INFRARED STEPTROSCOPY ON CARBON CHAIN MOLECULES: REVISED MEASUREMENTS ON C_{7}

\underline{THOMAS F. GIESEN}, PETRA NEUBAUER-GUENTHER, JÜRGEN KRIEG, STEPHAN SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln.

The formation of pure carbon chain molecules, C_{n} (n=3, 4, 5, ...), plays an important role in interstellar chemistry and in combustion processes as well. Spectroscopic studies and high level ab initio calculations have been performed in recent years to derive molecular properties such as structure, vibrational dynamics, and electronic configuration of pure carbon clusters. Ro-vibrational transitions of asymmetric stretching modes lie in the mid-infrared region at 5 μm whereas transitions from energetically low lying bending modes are expected to occur in the Terahertz region. To date only the bending mode of C_{3} has been directly measured by high resolution spectroscopy [1]. The search for bending mode transitions of longer carbon chain molecules is supported by recent large-scale coupled cluster calculations (see e.g. [2]) and by precise infrared measurements of asymmetric stretching modes and associated hot bands. Heath et al. used infrared laser absorption spectroscopy to measure the ν_{4} fundamental and associated ν_{11} hot band of C_{7} at 2137 cm^{−1} [3]. The authors found strong evidence for extremely large amplitude, anharmonic bending modes and concluded C_{7} to be a fairly floppy molecule. This finding is in contradiction to recent high level ab initio calculations by Botschwina [4] who found no evidence of floppiness for the C_{7} chain molecule, and therefore revised measurements are strongly needed. In this paper we present high resolution tunable infrared diode laser measurements on the C_{7} fundamental stretching mode ν_{4} and associated hot band transitions ν_{11} − (ν_{4} + ν_{11}) at 2137 cm^{−1} to clarify the question of C_{7} floppiness. Improved molecular constants for the fundamental band as well as for the hot band have been derived. We could not find any evidence for a large amplitude bending mode of C_{7}. All results are in very good agreement with calculations by Botschwina.

[1] C.A. Schmuttenmaer, R.C. Cohen, N. Pugliano et al., Science 249, 897 – 900 (1990).

[2] Botschwina Theoretical Chemistry Accounts 114, 350-356 (2005)

[3] J.R. Heath, R.J. Saykally, J. Chem. Phys. 94, 1724 (1991)

[4] P. Botschwina, Chem. Phys. Lett. 354, 148 (2002)